OpenStudy (anonymous):

Graph the line. x+4y=−8

OpenStudy (anonymous):

need help?

OpenStudy (anonymous):

yes please

OpenStudy (studygurl14):

www.desmos.com/calculator it is a graphing calculator

OpenStudy (anonymous):

so u need to graph it on a line graph or just get the answer

OpenStudy (anonymous):



OpenStudy (anonymous):

x-4y=-8 Start with the given equation -4y=-8-x Subtract x from both sides. -4y=-x-8 Rearrange the terms y=-x-8/-4 Divide both sides by -4 to isolate y. y=(-1/-4)x+-8/-4 Break up the fraction y=(1/4)x+2 Reduce. (0,2) is your graphing points

OpenStudy (anonymous):

i hope that helps